OpenStudy (anonymous):

Slope 0 and y intercept (0,6) equation is y=?

OpenStudy (anonymous):

Well, do you know in the equation of a line \[y=m(x-h)+k\] what each letter means?

OpenStudy (anonymous):

its been a long time

OpenStudy (anonymous):

Well, let's try to figure it out. How about this. Given two points, what's the slope of the line they make?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

im wrong

OpenStudy (anonymous):

No... that's an equation for the line. How about you tell me this. If you have \[p_1=(x, y)\] and \[p_2=(h, k)\], then what is the slope s you get between them?

OpenStudy (anonymous):

she just gave the slope, so that's a line

OpenStudy (anonymous):

0

OpenStudy (anonymous):

all she has to do is plug it in the slope-intercept formula

OpenStudy (anonymous):

Hmm... nvm. I was trying to teach, @panlac01 . Plug-n-chug is easy.

OpenStudy (anonymous):

im working on a test which deadline is by 11pm pst and im desperate for any help i can get

OpenStudy (anonymous):

listen to VF

OpenStudy (anonymous):

but you almost nailed it, candrie :)

OpenStudy (anonymous):

Okay, @candrie , when you're at the y-intercept, what's the coordinate pair?

OpenStudy (anonymous):

math is tough for me but i am awesome in english and everything else

OpenStudy (anonymous):

Oh the problem gives it to you. What is the y-intercept pair?

OpenStudy (anonymous):

0,6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you agree that the y-intercept is on the line, @candrie ?

OpenStudy (anonymous):

yes

OpenStudy (precal):

 in the end, remember to connect your final solution to its graph

OpenStudy (anonymous):

If a point is on a line, then its x and y values, when inserted into the formula you gave,\[y = m x +b\], must hold true.

OpenStudy (anonymous):

So put in the y-intercept's values into the formula.

OpenStudy (anonymous):

y=0(0)+6

OpenStudy (anonymous):

Well, I was trying to get you to say: 6=m(0)+b, so b = 6. But since you know what b and m are already, what's your question?

OpenStudy (anonymous):

just needed the equation

OpenStudy (anonymous):

Well it is y = m x + b. The difference is that my equation is y = m (x - h) + k But they're really the same because y = m x - mh + k y = m x +(k-mh) and k-mh =b

OpenStudy (anonymous):

y=mx+b y=0x+6 y=_____

OpenStudy (anonymous):

y=x+6?

OpenStudy (anonymous):

Multiply by 0...

OpenStudy (anonymous):

anything multiplied to zero is _____

OpenStudy (anonymous):

0

OpenStudy (anonymous):

therefore y=0x+6 (zero multiplied by x, plus six) y=_____

OpenStudy (anonymous):

the final answer is your equation, the whole thing

OpenStudy (anonymous):

y=6

OpenStudy (anonymous):

o wait

OpenStudy (anonymous):

y=0x+6?

OpenStudy (anonymous):

you got it already, but 0x can be further simplified by taking it away of the universe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

out of the universe

OpenStudy (anonymous):

my answer will be y=0+6?

OpenStudy (precal):

or y=6

OpenStudy (precal):